OpenStudy (anonymous):

Suppose that f(x)=((5x)/((2-2x)^5) Find an equation for the tangent line to the graph of at x=2.

OpenStudy (anonymous):

y= -5/16 x= 2 y'=((5(4x+1))/((32(x-1)^6))) y'(2)=45/32 45/32=slope (y-(-5/16))=45/32(x-2) (y+5/16)=45/32(x-2) (y+5/16)=45/32x-19/32 y=45/32x-29/32 this is what I came up with and it's wrong... Can someone help me?

OpenStudy (eyust707):

This step is incorrect: (y+5/16)=45/32(x-2) (y+5/16)=45/32x-19/32 

OpenStudy (eyust707):